OpenStudy (anonymous):

For the curve sqrt(xy)-x+y = 1, find the value of dy/dx at the point (9,4)

OpenStudy (dumbcow):

differentiate \[\frac{y + xy'}{2 \sqrt{xy}} -1 + y' = 0\] solve for y'

OpenStudy (anonymous):

but where did the 2sqrt(xy) come from?

OpenStudy (dumbcow):

it comes from derivative of sqrt(xy) you use chain rule \[\frac{d}{dx} \sqrt{u} = \frac{1}{2 \sqrt{u}} \frac{du}{dx}\]

OpenStudy (anonymous):

ok so for every square root, one must follow that correct?

OpenStudy (dumbcow):

yes it comes from power rule \[\sqrt{x} = x^{1/2} \rightarrow (x^{1/2})' = (1/2) x^{-1/2}\]

OpenStudy (anonymous):

oooh i see...

OpenStudy (anonymous):

now on the top part, the y + xy', is what the derivative of?

OpenStudy (dumbcow):

the "xy" inside the sqrt it comes from product rule

OpenStudy (anonymous):

oh i get it! now with that do we just plug in (9,4) into the x and y?

OpenStudy (dumbcow):

not yet, first you need to isolate y', then plug in the point to get value for y'

OpenStudy (dumbcow):

or i guess it doesn't matter, you can plug in the values first , then solve for y'

OpenStudy (anonymous):

allrighty!

OpenStudy (anonymous):

so does 3y'/4 = -3?

OpenStudy (anonymous):

i keep getting y' = 4/3, but that is not an answer choice, am i doing something wrong?

OpenStudy (dumbcow):

thats not what im getting

OpenStudy (dumbcow):

\[\frac{4 +9y'}{12} + y' = 1\] \[4 +9y' +12y' = 12\] \[21y' = 8\] ...

OpenStudy (anonymous):

wait, why did you make it =12 and have 12y'?

OpenStudy (dumbcow):

oh to get rid of the fraction, i multiplied everything by the denominator

OpenStudy (anonymous):

whoah! i didn't know you could do that! sudden realization of greater truth! then y' = 8/21

OpenStudy (dumbcow):

haha oh yeah, remember to do that it will make the algebra a bit easier :)

OpenStudy (anonymous):

allrighty! thank you sooo much!!

OpenStudy (dumbcow):

yw